Write the fraction as a percent.
1 1/5 (One and one/fifth)
can someone please help me with this? I have never been good with these kind of problems and obviously nothing has changed. I just do not understand how to do this.

Answer by BrittanyM(80)

In order to find the percentage equivilent of 1 1/5, we have to think of it as 6/5. (We get this by multiplying the 1 by 5, and adding it to the 1/5)
So, now that we have a single fraction, we can devide the numorator by the denomonator in order to get the number in the form of a decimal.
6/5 = 1.2 = 120%

We know that this is true because 5/5 is one which is equal to 100% and 1/5 is 20%.
Though this is the fastest way to solve the problem, you may find that it makes more sense to set up a proportion.
6/5 = x/100
600 = 5x
x=120%
**I'm in the process of writting a lesson plan on fractions and percentages. Many of the students on the site have been having problems with questions like this lately. I'm hoping to have it posted at the beginning of next week, if not sooner.